OpenStudy (anonymous):

Select all of the following that are ordered pairs of the given function. f(x) = 3 - 2x (-2, -1) (-1, 5) (0, 3) (1, 0) (2, -1)

OpenStudy (anonymous):

What ones do you think are correct? Just plug them into the equation and see if it is true or no

OpenStudy (anonymous):

i have no clue how to do that i mean i no where to put them but i dont no how to see if the answer is correct or not because i do not no how to solve them

OpenStudy (anonymous):

f(x) = 3 - 2x Lets try this point (-2, -1) -1 = 3 - 2(-2) -1 = 3 - 2(-2) -1 = 3 + 4 -1 = 7 Since -1 = 7 it is false and is not a solution.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

I took the equation f(x) = 3 - 2x and then I used the point, which is the first point you have (-2, -1) -2 is in the x position and -1 is in the y position. f() = y position and x where you see 2x = x position. take the x and y of the point (-2,-1) and plug it into your equation and solve like I did but do it with the rest of the points.

OpenStudy (anonymous):

oh okay